OpenStudy (anonymous):

What is the sum of a 6-term geometric sequence if the first term is 11, the last term is –11,264 and the common ratio is –4?

OpenStudy (zarkon):

-9009

OpenStudy (anonymous):

I'm still calculating. I'm a little slow, but I like to check my final answer with the pros :)

OpenStudy (a_clan):

a1[r^n -1] Sn= ------- ......Sum of sequence r-1 last term, b=a1(r^(n-1))

OpenStudy (anonymous):

OK. I also got negative 9009 :) I used A.clan formula so it takes me a while